# LaTeXTrans: Structured LaTeX Translation with Multi-Agent Coordination

### v2

@@ 5 Conclusion
Any instruction-following LLM can be integrated into our LaTeXTrans system. However, due to the large number of available models, it is impractical to evaluate each one individually. Therefore, we select a representative subset of commonly used LLMs for our experiments. We believe this selection sufficiently demonstrates the practicality and effectiveness of LaTeXTrans for LaTeX document translation. Additionally, although commercial systems such as Baidu and Youdao offer LaTeX translation services, they are not open-source. As a result, we are unable to compute metrics like COMETkiwi and FC-score for these systems. Therefore, we do not include a comprehensive comparison with them in our main experiments.
@@ Appendix A Additional Detailed Settings of the Experiment
When computing COMETkiwi and LLM-scores, we used pylatexenc‡‡‡https://github.com/phfaist/pylatexenc to convert each LaTeX translation unit into plain text. Although LaTeXTrans parses structured LaTeX documents into fine-grained translation units, we followed the baseline’s evaluation protocol by using section-level translation units for computing both COMETkiwi and LLM-scores. Furthermore, to assess contextual consistency in the LLM-score evaluation, we concatenated section-level translation units into paired paragraphs and then scored them using GPT-4o. The prompt template used for scoring is illustrated in Figure 12. When calculating the FC-score, we set the initial score to 100. Since errors have a greater impact on the final PDF format scheduling effect than warnings, in the experiment, we set the value of (10) to be significantly greater than (2). Ultimately, whether the compilation is successful is the most intuitive factor for evaluating the compilation. Therefore, in the experiment, we set the to 20.
Datasets
We selected the LaTeX source files of 50 academic papers in the field of computer science from arXiv as our test set. The distribution of paper lengths is shown in Figure 4. Additionally, we analyzed the topics of the papers and visualized them as a word cloud in Figure 5. This result shows that the test set exhibits a diverse range of paper lengths, covering both short and long documents, which helps ensure robustness across different document sizes. Moreover, the word cloud reveals a wide variety of research topics within the computer science domain, confirming the topical diversity of the test set and enhancing the generality of our evaluation.
Appendix B System Performance Display
We select six cases to visually demonstrate the translation performance of our system, focusing on En-Zh and En-Ja translation tasks, as illustrated in Figure 6 to Figure 11. All six cases are translation cases of the LaTeX source code of papers by LaTeXTrans. In each case, we have selected two relatively complex parts to present. Among the six cases, there are three En-Zh translation tasks and three En-Ja translation tasks, respectively.
Appendix C Prompt Templates for LLM-Based Components in LaTeXTrans

### v3

@@ 6 Conclusion
Any instruction-following LLM can be integrated into our LaTeXTrans system. However, due to the large number of available models, it is impractical to evaluate each one individually. Therefore, we select a representative subset of commonly used LLMs for our experiments. We believe this selection sufficiently demonstrates the practicality and effectiveness of LaTeXTrans for LaTeX document translation. Additionally, although commercial systems such as Baidu and Youdao offer LaTeX translation services, they are not open-source. As a result, we are unable to compute metrics like COMETkiwi and LLM-score for these systems. Therefore, we do not include a comprehensive comparison with them in our main experiments.
@@ Appendix A Additional Detailed Settings of the Experiment
We use GPT-4o to calculate LLM-scores. When computing COMETkiwi and LLM-scores, we used pylatexenc******https://github.com/phfaist/pylatexenc to convert each LaTeX translation unit into plain text. Although LaTeXTrans parses structured LaTeX documents into fine-grained translation units, we followed the baseline’s evaluation protocol by using section-level translation units for computing both COMETkiwi and LLM-scores. Furthermore, to assess contextual consistency in the LLM-score evaluation, we concatenated section-level translation units into paired paragraphs and then scored them using GPT-4o.
Appendix B Formatting Error Categories
We categorize formatting-related errors into two major types:
Structure-induced formatting errors, arising from corrupted or improperly restored LaTeX commands. These may result in missing formatting effects (e.g., bold text not rendered), incorrect cross-references, or garbled content in the compiled document.
Content-level errors, involving incorrect or unintended textual outputs, such as untranslated segments or irrelevant words.
Figure 4 illustrates representative examples of formatting errors.
To ensure consistent evaluation, we adopt the following counting rules. For structure-induced errors, if a single corrupted LaTeX command causes multiple downstream anomalies (e.g., numerous missing cross-references triggered by one faulty reference definition), the error is counted only once per document. In contrast, content-level errors are counted independently at the instance level. This design prevents inflated error counts caused by cascading structural failures while preserving fine-grained sensitivity to semantic and textual inaccuracies.
Appendix C Case Study
We present two case studies to visually demonstrate the translation performance of LaTeXTrans on the En–Zh and En–Ja tasks, as shown in Figure 5. Each case corresponds to the translation of a real-world LaTeX source document. For illustration, we highlight two relatively complex segments in each example.